OpenStudy (anonymous):

how do i solve x^4-2x^3=16+8x-4x^3

OpenStudy (anonymous):

okay. first you need to being all the terms to one side. So bring in the 16+8x - 4x^3 to the left hand side of the equation and post what you got here.

OpenStudy (anonymous):

x^4+2x^3-8x-16=0

OpenStudy (anonymous):

okay great. What is common in the first two terms x^4 and 2x^3?

OpenStudy (anonymous):

x^2

OpenStudy (anonymous):

are you sure about that?

OpenStudy (anonymous):

no its x^3

OpenStudy (anonymous):

okay. so factor out the x^3 in the first two terms and post what you got.

OpenStudy (anonymous):

x^3(x+2)

OpenStudy (anonymous):

good! next lets move on to -8x -16. what is common in those two terms?

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

what is 8 times 2?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

so its 8..

OpenStudy (anonymous):

okay what is -8 times 2?

OpenStudy (anonymous):

-16

OpenStudy (anonymous):

so what can you factor out?

OpenStudy (anonymous):

im confused..

OpenStudy (anonymous):

-8x-16. what can you factor out here?

OpenStudy (anonymous):

-8(x-2)

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

no -8(x+2)

OpenStudy (anonymous):

yes, that is correct.

OpenStudy (anonymous):

-8(x+2) is correct

OpenStudy (anonymous):

so the whole equation is what X^3(x+2)-8(x+2) right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

im sorry, it is x^3(x+2)-8(x+2)=0

OpenStudy (anonymous):

so what can you factor out here?

OpenStudy (anonymous):

(x^3-8)(x+2)=0 x=-2,2 is the answer

OpenStudy (anonymous):

thinker, I was in the middle of explaining the whole problem from beginning to end so that car could do it on his own without me thinking for him and doing his calculations for him.

OpenStudy (anonymous):

i wanted to figure it out myself

OpenStudy (anonymous):

that is not how he will learn.

OpenStudy (anonymous):

ohh...sorry..just wanted to help :)

OpenStudy (anonymous):

I understand :)

OpenStudy (anonymous):

thanks though guys!

OpenStudy (anonymous):

youre welcome.

OpenStudy (anonymous):

ur welcome ! :)

OpenStudy (anonymous):

by the way im a girl lol

OpenStudy (anonymous):

sorry, lol

OpenStudy (anonymous):

alright car, good luck .Look for patterns when you are asked to solve or simplify equations.

OpenStudy (anonymous):

ok!

OpenStudy (anonymous):

what if the problem is x/x^2-4 + 1/x+2 =3..do i solve it the same by bringing the 3 to the other side?